OpenStudy (meilendurcer):

How does finding the square root [sqrt] of a number compare to finding the cube root [description] of a number? Use the number 64 in your explanation.

OpenStudy (meilendurcer):

[\[\sqrt[3]{?}\]

OpenStudy (anonymous):

for the square root, you want the number whose square is inside the radial, so for example the square root of 64 is 8 because \(8^2=64\)

OpenStudy (anonymous):

for the cubed root, you want the number whose CUBE is inside the radical for example if the number is 64 you are solving \(a^3=64\) and since \(4^3=64\) you have \(\sqrt[3]{64}=4\)